# PhysGM: Large Physical Gaussian Model for Feed-Forward 4D Synthesis

### v1

@@ Conclusion
Method
No Pre-opt.
3D Gaussians
Auto Param
Computation
Generalizable
Without LLM
Inference Time
PhysGaussian (Xie et al. 2024) $\times$ $\times$ $\times$ $\checkmark$ -
DreamPhysics (Huang et al. 2025) $\times$ only E $\times$ $\checkmark$ >0.5h
PhysDreamer (Zhang et al. 2024c) $\times$ only E $\times$ $\checkmark$ >1h
OmniPhysGS (Lin et al. 2025) $\times$ only material $\times$ $\checkmark$ >12h
DreamGaussian4D (Ren et al. 2023) $\checkmark$ $\times$ $\checkmark$ $\times$ 6.5min
Feature Splatting (Qiu et al. 2024) $\times$ $\times$ $\times$ $\checkmark$ >1h
PhysSplat (Zhao et al. 2024) $\times$ $\checkmark$ $\checkmark$ $\times$ <2min
@@
More Details on Implementation
Compare with Other Methods
Table 2 provides a qualitative comparison between our method, PhysGM, and other state-of-the-art approaches (Xie et al. 2024; Zhang et al. 2024c; Huang et al. 2025; Lin et al. 2025; Ren et al. 2023; Qiu et al. 2024; Zhao et al. 2024). We evaluate each method across five critical dimensions: two concerning input requirements (the need for pre-optimized 3D Gaussians or pre-defined physical parameter) and three concerning core capabilities (generalizability, independence from Large Language Models, and inference speed). The comparison highlights that our approach is the only one to operate without these stringent prerequisites. PhysGM simultaneously achieves strong generalization and maintains a very short inference time of under 30 seconds.
Simulation Details
@@
MPM Grid Resolution: The simulation domain is discretized into a background grid of 200 * 200 * 200 cells. This grid is fundamental to the MPM algorithm for computing particle interactions and mapping data between particles and the grid.
Camera Motion: The camera is configured to be static during the simulation.
@@
Other Parameters: Gravity is applied in the falling scene, and force in the corresponding direction is applied in the collision scene.
Training Details
We train our model on 8 NVIDIA A800 GPUs using a two-stage process for 12 hours in total. The stage one consisted of 30 epochs, and the stage two consisted of 20 epochs. We employed the AdamW optimizer with a learning rate of 1e-6 and applied gradient clipping with a maximum norm of 1.0. For data augmentation, both grid distortion and camera jitter were applied with a probability of 50%.
Material Constitutive Models
@@
The Fixed Corotational Constitutive Model
@@
The Drucker-Prager Plasticity Model
@@
Dataset
This section provides a detailed account of the construction process for our PhysAssets Dataset, which serves as the foundation for our research. The final dataset comprises over 24,000 unique 3D objects, meticulously curated and integrated from four prominent existing datasets: OmniObject3D (Wu et al. 2023), Google Scanned Objects (GSO) (Downs et al. 2022), Amazon Berkeley Objects (ABO) (Collins et al. 2022), and Objaverse (Deitke et al. 2023). The primary objective of this effort was to create a comprehensive, diverse, and standardized collection of Physical-based assets annotated with four-view rendered images, physical properties, and corresponding guiding videos.
Source Datasets
@@
Google Scanned Objects (GSO)
This dataset contains 1,030 photorealistic 3D models of common household items, such as kitchenware, stationery, and electronics. The models are provided as textured meshes (.obj, .mtl) and include pre-rendered orthogonal views (front, back, left, right), which facilitate standardized data extraction for our pipeline.
@@
To ensure quality and consistency across the heterogeneous source datasets, we established a systematic data curation and processing pipeline. The Objaverse dataset, while extensive, is characterized by its considerable size and variable data quality. Consequently, to extract a high-quality subset, we employed a systematic curation strategy analogous to the one applied to gobjaverse. The screening procedure is outlined as follows: (1) A geometric similarity clustering algorithm was employed to identify and remove near-duplicate models. Any model exhibiting a similarity score of over 85% with another was considered redundant and removed; (2) To filter out objects with non-standard or incomplete textures, we performed an analysis in the HSV (Hue, Saturation, Value) color space. Models where white pixels constituted more than 75% of the surface texture were discarded, as this often indicates missing or placeholder textures. In the end, we filtered approximately 10k data points in the Objaverse.For the other datasets, we used the full data without applying a filtering process. For datasets that do not provide a four view rendering view, we use the rendering script provided by the corresponding dataset for four view rendering. This procedure ensures that every object in our dataset is represented by a consistent set of views, capturing its complete geometric features for subsequent learning tasks.
Physical Property Annotation
We employ a vision-language joint modeling approach for material classification. The four orthogonal RGB images of each object are fed into a pre-trained SigLIP-2 model. This model calculates the cross-modal similarity between the object’s visual features and the textual descriptions of seventeen common materials (e.g., “wood”, “metal”, “plasticine”). The final material label for each object is determined via a weighted voting mechanism based on these similarity scores. The prediction of mechanical parameters, specifically Young’s modulus and the Poisson’s ratio, is achieved through a similar methodology. Once a material is identified, its density value (in g/cm³) is assigned using a layered mapping framework that correlates the material label with standard physical constants.
Video Generation and Preference Calculation Process
@@
To facilitate the second stage of our training, which employs Direct Preference Optimization (DPO), we established a systematic pipeline for generating a large-scale dataset of preference tuples. This process is crucial for providing the high-quality, ranked data required to fine-tune our model on the nuances of physical dynamics. The pipeline consists of three main steps:
Ground-Truth Videos Generation
For some scenes in our dataset, we generate a high-fidelity, ground-truth simulation video. This is achieved by utilizing FramePack, a video generation model, which is guided by a unique, descriptive prompt for each material and then manually selected. These prompts define the initial conditions and interactions, ensuring the resulting video serves as an authoritative benchmark for both physical accuracy and visual realism.
Candidate Video Generation

### v2

@@ 6 Conclusion
Method No Pre-opt. 3D Gaussians Auto Param Computation Generalizable Without LLM Inference Time
PhysGaussian [xie2024physgaussian] $\times$ $\times$ $\times$ $\checkmark$ -
DreamPhysics [huang2025dreamphysics] $\times$ only E $\times$ $\checkmark$ >0.5h
PhysDreamer [zhang2024physdreamer] $\times$ only E $\times$ $\checkmark$ >1h
OmniPhysGS [lin2025omniphysgs] $\times$ only material $\times$ $\checkmark$ >12h
DreamGaussian4D [ren2023dreamgaussian4d] $\checkmark$ $\times$ $\checkmark$ $\times$ 6.5min
Feature Splatting [qiu2024feature] $\times$ $\times$ $\times$ $\checkmark$ >1h
PhysSplat [zhao2024efficient] $\times$ $\checkmark$ $\checkmark$ $\times$ <2min
@@
Appendix A More Details on Implementation
A.1 Compare with Other Methods
Table 4 provides a qualitative comparison between our method, PhysGM, and other state-of-the-art approaches [xie2024physgaussian, zhang2024physdreamer, huang2025dreamphysics, lin2025omniphysgs, ren2023dreamgaussian4d, qiu2024feature, zhao2024efficient]. We evaluate each method across five critical dimensions: two concerning input requirements (the need for pre-optimized 3D Gaussians or pre-defined physical parameter) and three concerning core capabilities (generalizability, independence from Large Language Models, and inference speed). The comparison highlights that our approach is the only one to operate without these stringent prerequisites. PhysGM simultaneously achieves strong generalization and maintains a very short inference time of under 30 seconds.
A.2 Simulation Details
@@
The simulation domain is discretized into a background grid of 200 * 200 * 200 cells. This grid is fundamental to the MPM algorithm for computing particle interactions and mapping data between particles and the grid.
Camera Position
For different objects, the camera is initialized at an azimuth of -45 or 135 degrees, an elevation of 0 degrees, and a radius of 1.8 or 1.3 units.
Camera Motion
The camera is configured to be static during the simulation.
Other Parameters
@@
Gravity is applied in the falling scene, and force in the corresponding direction is applied in the collision scene.
A.3 Training and Evaluation Details
Network Architecture.
We employ DINOv3 [siméoni2025dinov3] (ViT-L/16) pre-trained on LVD-1689M as our image encoder, producing 1024-dimensional features. The transformer backbone consists of 24 layers with a hidden dimension of 1024 and attention head dimension of 64. We use a patch size of 16 and incorporate 3 learnable global tokens for physics prediction. For 3D Gaussian representation, we set the spherical harmonics degree to 0, with near and far planes at 0.001 and 2.0, respectively.
Training Configuration.
We train our model on 32 NVIDIA A800 GPUs using a two-stage process for about 3 days in total with a batch size of 8 per GPU. The base learning rate is set to $2*10^{-4}$ with AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay == 0.05). We employ a cosine learning rate schedule with 5K warmup steps and clip gradients to a maximum norm of 10.0. Mixed precision training is enabled using bfloat16 to accelerate computation.
Data Configuration.
During training, input images are resized to $512\times 512$ resolution with square cropping. Each training sample consists of 4 input views and 8 target views, where target views include the input views for consistency. We use 8 workers for data loading with a prefetch factor of 128 to ensure efficient GPU utilization.
Evaluation Protocol.
We evaluate on the complete GSO dataset [downs2022google] containing 1,009 objects. For each object, we render 32 views with 4 elevation angles ($0\textdegree,20\textdegree,40\textdegree,60\textdegree$) and 8 azimuthal angles. During testing, we sample fixed 4 views as input and evaluate reconstruction quality on 8 randomly selected novel views. We report PSNR, SSIM [wang2004image], and LPIPS [zhang2018unreasonable] averaged across all test views and objects.
A.4 User Preference Evaluation
To complement quantitative metrics with human perception assessment, we conduct a user study to evaluate the perceptual quality and physical plausibility of generated 4D sequences across different methods.
Study Design.
We employ a Four-Alternative Forced Choice (4AFC) protocol, where participants are presented with four videos simultaneously showing the same object simulated by different methods: our PhysGM, and three baseline methods (PhysGaussian [xie2024physgaussian], OmniPhysGS [lin2025omniphysgs], and DreamGaussian4D [ren2023dreamgaussian4d]). The videos are displayed in randomized positions to eliminate order bias. Participants are instructed to select the single video that exhibits the most realistic physical behavior and visual quality, considering factors such as motion naturalness, material response, temporal coherence, and rendering fidelity.
Stimuli and Sampling.
We carefully select 5 representative test scenes spanning diverse object categories and physical scenarios (dropping and stretching). For each scene, we generate 4D sequences using all methods with identical input views and physical interaction setups to ensure fair comparison.
Participants and Procedure.
We recruited 103 participants comprising graduate students and researchers with backgrounds in computer graphics, computer vision, or related fields. Each participant completed a questionnaire containing 5 comparison trials (one per test scene). Before the formal study, participants underwent a training phase with two practice trials to familiarize themselves with the task and interface. Participants could replay videos multiple times before making their selection and were allowed to take breaks between trials. The entire study took approximately 10 minutes per participant.
Data Validation and Filtering.
To ensure data quality, we implemented several validation mechanisms:
  • Attention checks: Two control trials with obvious quality differences were inserted to identify inattentive participants.
  • Completion time: Responses completed too quickly (<<5 seconds per trial) were flagged.
  • Response consistency: Participants showing random selection patterns were identified via entropy analysis.
After applying these criteria, we excluded 3 invalid responses (2.9% exclusion rate) due to failed attention checks or suspiciously short completion times, resulting in 100 valid responses for analysis.
User Preference Rate (UPR).
We define the User Preference Rate as the percentage of participants who selected a given method as the most realistic:
$\text{UPR}_{m}=\frac{1}{S\cdot N}\sum_{s=1}^{S}\sum_{i=1}^{N}\mathbf{1}[\text{choice}_{s,i}=m]\times 100\%$ (11)
where $m$ denotes the method, $S$ is the number of test scenes, $N=100$ is the number of valid participants, and $\mathbf{1}[\cdot]$ is the indicator function. A higher UPR indicates stronger human preference. Under random chance, each method would receive 25% preference rate in a 4AFC setup.
Appendix B Material Constitutive Models
@@ B.1 The Neo-Hookean Model
B.2 The Fixed Corotational Constitutive Model
@@
B.3 The Drucker-Prager Plasticity Model
@@
Table 5: Material distribution in PhysAssets dataset. The 14 primary materials account for 97% of the dataset, while 32 rare materials provide additional diversity.
Rank Material Count Percentage
1 Plastic 13,696 27.3%
2 Wood 8,443 16.8%
3 Metal 7,353 14.6%
4 Fabric 7,255 14.5%
5 Ceramic 3,023 6.0%
6 Stone 2,135 4.3%
7 Paper 1,432 2.9%
8 Leather 1,132 2.3%
9 Glass 955 1.9%
10 Rubber 687 1.4%
11 Foam 168 0.3%
12 Snow 147 0.3%
13 Sand 58 0.1%
14 Other (32 materials) 1,731 3.4%
Appendix C PhysAssets Dataset Statistics
C.1 Dataset Composition
PhysAssets comprises a comprehensive collection of 3D assets with annotated physical properties. The dataset consists of two main components: a training set containing 49,206 objects aggregated from multiple public repositories (Objaverse [deitke2023objaverse], OmniObject3D [wu2023omniobject3d], ABO [collins2022abo], and HSSD [khanna2023hssd]), and a held-out test set of 1,009 objects from the Google Scanned Objects (GSO) dataset [downs2022google], totaling 50,215 annotated 3D objects. The primary objective of this effort was to create a comprehensive, diverse, and standardized collection of Physical-based assets annotated with 20+ views rendered images, physical properties, and corresponding guiding videos.
C.2 Material and Physical Property Distribution
The dataset exhibits rich material diversity, covering 46 distinct material categories. Among these, 14 primary materials constitute the majority of the dataset, while 32 additional rare materials provide coverage for specialized physical scenarios. Table 5 presents the distribution of the 14 primary materials.
The most represented material is Plastic, with 13,696 samples (27.3%), reflecting its prevalence in manufactured objects. Wood constitutes the second largest category with 8,443 samples (16.8%), followed by Metal with 7,353 samples (14.6%) and Fabric with 7,255 samples (14.5%). Ceramic objects account for 3,023 samples (6.0%). Medium-frequency materials include Stone (2,135 samples, 4.3%), Paper (1,432 samples, 2.9%), Leather (1,132 samples, 2.3%), Glass (955 samples, 1.9%), and Rubber (687 samples, 1.4%). Low-frequency but physically interesting materials comprise Foam (168 samples, 0.3%), Snow (147 samples, 0.3%), and Sand (58 samples, 0.1%). The remaining 32 rare materials collectively account for approximately 3.0% of the dataset, providing diversity for edge cases and specialized physical behaviors.
This heterogeneous material distribution enables our model to learn a comprehensive physical prior spanning rigid bodies (metal, stone), deformable materials (rubber, foam), granular substances (sand, snow), and everyday materials (plastic, wood, fabric). The long-tail distribution also facilitates studying generalization to rare material types. Young’s Modulus ($E$): Measures material stiffness, ranging from soft materials ($10^{3}$ Pa) to rigid materials ($4*10^{11}$ Pa). The dataset contains 10 distinct values spanning this range.
Poisson’s Ratio ($\nu$): Characterizes material compressibility, typically ranging from 0.01 to 0.49. The dataset includes 10 representative values covering common material behaviors.
C.3 Source Datasets
@@
HSSD Dataset.
The Habitat Synthetic Scenes Dataset (HSSD) [khanna2023hssd], contains over 18,000 high-quality indoor scenes with photorealistic rendering and detailed semantic annotations. The dataset features diverse residential and commercial environments with realistic layouts and furnishings.
@@ C.4 Data Processing
To ensure quality and consistency across the heterogeneous source datasets, we established a systematic data curation and processing pipeline. The Objaverse dataset, while extensive, is characterized by its considerable size and variable data quality. Consequently, to extract a high-quality subset, we employed a systematic curation strategy analogous to the one applied to gobjaverse. The screening procedure is outlined as follows: (1) A geometric similarity clustering algorithm was employed to identify and remove near-duplicate models. Any model exhibiting a similarity score of over 85% with another was considered redundant and removed; (2) To filter out objects with non-standard or incomplete textures, we performed an analysis in the HSV (Hue, Saturation, Value) color space. Models where white pixels constituted more than 75% of the surface texture were discarded, as this often indicates missing or placeholder textures. In the end, we filtered approximately about 20k data points in the Objaverse.For the other datasets, we used the full data without applying a filtering process. For datasets that do not provide enough view rendering view, we use the rendering script provided by the corresponding dataset for enough view rendering. This procedure ensures that every object in our dataset is represented by a consistent set of views, capturing its complete geometric features for subsequent learning tasks.
Figure 7: Automated dataset construction pipeline. We predict physical properties using Qwen3-VL, and generate reference videos using FramePack.
Appendix D Dataset Construction Pipeline.
As shown in Figure 7, we construct PhysAssets through an automated pipeline which predicting physical properties (material class, Young’s modulus, Poisson’s ratio) from 8 selected views using Qwen3-VL [qwen3technicalreport] and generating ground-truth reference videos using FramePack [zhang2025packing] conditioned on predicted properties. This pipeline enables scalable annotation of 50,215 objects with physical properties and reference dynamics.
D.1 Physical Property Annotation Pipeline
We develop a semi-automatic annotation pipeline leveraging multimodal large language models to predict three critical physical properties for each object: material class, Young’s modulus ($E$), and Poisson’s ratio ($\nu$). This approach enables scalable annotation of large-scale 3D datasets while maintaining consistency with real-world material physics.
D.1.1 Visual Feature Extraction
For each 3D object, we choose eight uniformly distributed views at fixed elevation. These multi-view RGB images provide comprehensive visual coverage of the object’s geometry, texture, and appearance. The views are then fed into Qwen3-VL [qwen3technicalreport], a state-of-the-art vision-language model pre-trained on diverse visual and textual data.
Material Classification
Material classification is performed through vision-language alignment. We define a closed vocabulary of 45 primary materials commonly found in everyday objects: Wood, Metal, Plastic, Glass, Fabric, Leather, Ceramic, Stone, Rubber, Paper, Sand, Snow, Plasticine, Foam, etc.. The model is queried with the following prompt:
Material Classification Prompt:
“What is the primary material of the object in these images? Answer with a single word from this list: Wood, Metal, Plastic, Glass, Fabric, Leather, Ceramic, Stone, Rubber, Paper, Sand, Snow, Plasticine, Foam, etc.”
The model processes all eight views and outputs a material label based on cross-modal similarity between visual features and material descriptions. In cases of ambiguity, a weighted voting mechanism across views determines the final material class.
Young’s Modulus Prediction
Young’s modulus ($E$) characterizes material stiffness—the resistance to elastic deformation under stress. We discretize the continuous range of Young’s modulus values into 10 interpretable categories spanning from extremely soft materials (e.g., gel, foam) to ultra-stiff materials (e.g., diamond, tungsten). The model is prompted with:
“Based on these images, determine the Young’s Modulus ($E$) of the object.
What is Young’s Modulus?
Young’s Modulus ($E$) measures a material’s stiffness or resistance to elastic deformation. It indicates how much stress is needed to produce a given amount of strain (deformation).
Select the most appropriate description:
  1. 1.
    extremely soft - Like gel or foam (e.g., jelly, soft foam) $\sim$ 1 KPa
  2. 2.
    very soft - Like rubber or sponge (e.g., rubber bands, foam mattress) $\sim$100 KPa
  3. 3.
    soft - Like soft plastics or leather (e.g., leather, soft PVC) $\sim$1 MPa
  4. 4.
    moderately soft - Like hard rubber (e.g., tire rubber) $\sim$10 MPa
  5. 5.
    moderate - Like nylon or cork (e.g., nylon, wood cork) $\sim$100 MPa
  6. 6.
    moderately stiff - Like hard plastics (e.g., ABS plastic, acrylic) $\sim$1 GPa
  7. 7.
    stiff - Like glass or ceramics (e.g., glass, porcelain) $\sim$10 GPa
  8. 8.
    very stiff - Like aluminum (e.g., aluminum, brass) $\sim$70 GPa
  9. 9.
    extremely stiff - Like steel (e.g., steel, iron) $\sim$200 GPa
  10. 10.
    ultra stiff - Like tungsten or diamond (e.g., tungsten, diamond) $\sim$400 GPa
Answer with ONLY ONE of these exact keywords.”
The predicted categorical label is then mapped to a numerical value in Pascals (Pa) using the mapping defined in Table 6.
Poisson’s Ratio Prediction
Poisson’s ratio ($\nu$) quantifies the ratio of lateral strain to axial strain when a material is deformed. We similarly discretize Poisson’s ratio into 10 categories representing different material behaviors, from auxetic materials (negative Poisson’s ratio) to nearly incompressible materials (approaching 0.5). The prediction prompt is:
Poisson’s Ratio Prediction Prompt:
“Based on these images, determine the Poisson’s Ratio ($\nu$) of the object.
What is Poisson’s Ratio?
Poisson’s Ratio ($\nu$) measures how much a material expands laterally when compressed axially, or contracts laterally when stretched. It describes the relationship between lateral strain and axial strain.
Select the most appropriate description:
  1. 1.
    nearly incompressible - Almost no volume change (e.g., rubber) $\sim$0.50
  2. 2.
    high resistance - High lateral expansion (e.g., soft rubber) $\sim$0.45
  3. 3.
    moderately high - Moderately high deformation (e.g., gold, lead) $\sim$0.40
  4. 4.
    moderate high - Above average deformation (e.g., plastic, aluminum) $\sim$0.35
  5. 5.
    moderate - Typical for many metals (e.g., steel, iron) $\sim$0.30
  6. 6.
    moderate low - Below average deformation (e.g., glass) $\sim$0.25
  7. 7.
    low - Low lateral expansion (e.g., concrete, ceramics) $\sim$0.20
  8. 8.
    very low - Very low lateral expansion (e.g., cork) $\sim$0.15
  9. 9.
    extremely low - Minimal lateral deformation (e.g., foam) $\sim$0.10
  10. 10.
    auxetic - Negative Poisson’s ratio materials $\sim$0.01
Answer with ONLY ONE of these exact keywords.”
The categorical output is converted to a dimensionless numerical value using the mapping in Table 7.
Property Mapping Tables
The categorical predictions from the vision-language model are mapped to numerical physical property values suitable for Material Point Method (MPM) simulation. Tables 6 and 7 present the complete mappings.
Table 6: Young’s Modulus categorical to numerical mapping. Values span 8 orders of magnitude, covering materials from soft gels to ultra-hard ceramics.
Category Example Materials Value (Pa)
extremely soft Gel, foam, jelly $1.0\times 10^{3}$
very soft Rubber, sponge, silicone $1.0\times 10^{5}$
soft Leather, soft PVC, fabric $1.0\times 10^{6}$
moderately soft Hard rubber, tire rubber $1.0\times 10^{7}$
moderate Nylon, cork, paper $1.0\times 10^{8}$
moderately stiff Hard plastic (ABS, acrylic) $1.0\times 10^{9}$
stiff Glass, ceramic, porcelain $1.0\times 10^{10}$
very stiff Aluminum, brass, bronze $7.0\times 10^{10}$
extremely stiff Steel, iron, stainless steel $2.0\times 10^{11}$
ultra stiff Tungsten, diamond, carbide $4.0\times 10^{11}$
Table 7: Poisson’s Ratio categorical to numerical mapping. Values range from 0.01 (auxetic materials) to 0.49 (nearly incompressible materials).
Category Example Materials Value
auxetic Special engineered materials 0.01
extremely low Foam materials 0.10
very low Cork, engineered materials 0.15
low Concrete, ceramics, brick 0.20
moderate low Glass, cast iron 0.25
moderate Steel, iron, brass, titanium 0.30
moderate high Plastic, aluminum, copper 0.35
moderately high Gold, lead, clay 0.40
high resistance Soft rubber, flexible polymers 0.45
nearly incompressible Rubber, elastomers 0.49
D.2 Video Generation and Preference Calculation
@@
To facilitate the second stage of our training, which employs Direct Preference Optimization (DPO), we established a systematic pipeline for generating a dataset of preference tuples. This process is crucial for providing the high-quality, ranked data required to fine-tune our model on the nuances of physical dynamics. The pipeline consists of three main steps:
D.2.1 Ground-Truth Video Generation
We generate reference videos using FramePack [zhang2025packing], guided by text prompts describing the physical scenario. After evaluating multiple prompt formulations (detailed below), we selected the following template for its optimal balance of simplicity and physical realism:
“An object made of {material} falls straight down from the air onto the invisible ground, white background, no extra background, no shadow.”
D.2.2 Alternative Prompt Variants
For reference and reproducibility, we document the alternative prompt variants explored during our experimentation. These prompts represent different trade-offs between prompt complexity, physical constraints, and generation control.
Prompt 2 (Detailed Physics Description):
“A {material} toy centered on a plain pure white background. The {material} toy falls straight down vertically from the center of the frame to the bottom edge, obeying the laws of physics (gravity, acceleration). Show the entire descent: starting stationary at center, accelerating downwards, hitting the bottom edge with a subtle impact, and coming to a complete stop. The {material} toy remains rigid and inanimate throughout, showing no deformation or independent movement. Fixed, static camera view. No anthropomorphism, no unexpected motion, only the physics-based vertical fall and stop.”
Limitation: Overly detailed constraints sometimes led to inconsistent generation or failure to satisfy all specified conditions.
Prompt 5 (Identity Preservation Focus):
“Generate a short, high-fidelity video based on the provided object image, where the absolute highest priority is to strictly maintain the object’s identity throughout the entire sequence. The scene features a seamless white background and a solid, invisible, horizontal white floor. The video begins with the object perfectly still in mid-air, then it is released to fall straight down vertically under gravity. Crucially, the object must maintain its initial orientation during the fall, without any tumbling, spinning, or rotation. The object is made of {material}, and its impact and subsequent behavior must realistically simulate the physical properties of this material.”
Limitation: While improving identity preservation, this prompt occasionally resulted in unrealistic motion due to strict orientation constraints.
Prompt 6 (Photorealism Emphasis):
“Generate a short, photorealistic video based on the provided input image, simulating the object falling and impacting the ground. Throughout the entire video, the object must retain its original visual identity—its shape, texture, and color. The fall itself must be completely inanimate and passive; the object must descend in a pure vertical drop without any rotation, spinning, or tumbling. Upon impact with the flat white ground, the object’s physical reaction must precisely mimic the properties of {material}. The entire event takes place in a seamless, infinite white studio environment.”
Limitation: We found that excessively long prompts with detailed constraints often compromise generation quality, leading to inconsistent or unnatural motion.
The selected prompt consistently produced the most physically plausible and visually coherent videos across diverse materials and object geometries.
D.2.3 Candidate Video Generation